# Spatial mixing of stellar populations in globular clusters via binary–single star scattering

### v1

@@ 1 Introduction
Star clusters are often used as laboratories for studying star formation, stellar evolution and the dynamics of galaxies. When focusing on the old ($2\,\mathrm{Gyr}$) and relatively massive ($10^{4}\,M_{\odot}$) globular clusters (GCs), these gravitationally bound systems have long been known to host multiple populations of stars with distinct chemical compositions and evolutionary histories (see, e.g. Carretta et al., 2010; Gratton et al., 2012; Bellini et al., 2015; Bastian & Lardo, 2018; Martocchia et al., 2018; Libralato et al., 2019; Tiongco et al., 2019; Martocchia, 2020; Sollima, 2021; Vesperini et al., 2021, and references therein). Hereafter, stars from the primordial (or first) population are referred to as ‘P1’, and the enriched second population as ‘P2’. For instance, variations in the abundances of light elements point to complex formation scenarios of P1 and P2, including the accretion of enriched material onto existing stars or a direct formation of P2 stars out of this enriched material (e.g. Carretta et al., 2010; Bekki, 2011). In some clusters, P1 and P2 are also characterised by different dynamical evolution, which leads to measurable differences in present-day kinematic properties, such as rotation or anisotropy in velocity dispersion (Cordoni et al., 2020b, a; Dalessandro et al., 2024; Leitinger et al., 2025; Cadelano et al., 2024). Since velocity anisotropy affects the relaxation time scales, mass segregation, and the evolution towards energy equipartition (see, e.g. Tiongco et al., 2016; Bianchini et al., 2017; Pavl´ık & Vesperini, 2021, 2022a, 2022b; Pavl´ık et al., 2024), it is expected that each stellar population might evolve differently (as shown, e.g. by Livernois et al., 2024).
@@
Most models of GC formation predict that P2 stars are being preferentially born in their cores (e.g. Lacchin et al., 2022; Yaghoobi et al., 2022a, b). Although the spatial distribution of these populations evolves due to relaxation processes or gas expulsion (e.g. Decressin et al., 2010), dynamically younger GCs are expected to retain their initial structural differences (e.g. Dalessandro et al., 2019, 2024, and references therein). However, a recent observational study by Leitinger et al. (2023) showed that the spatial distribution of P1 and P2 varies among the dynamically-young Galactic GCs. Notably, while P2 stars in some GCs (e.g. NGC 5024 or NGC 6809) appear to be more centrally concentrated than P1 (as expected for GCs less than a few relaxation times old), in other clusters which have similar dynamical ages as these ones (e.g. NGC 3201 or NGC 6101), P2 stars primarily occupy the outer regions of the clusters and P1 is preferentially in the central regions. There are also several dynamically-young GCs which are fully mixed, i.e. they show no preferred concentration of either P1 or P2 (e.g. NGC 288, NGC 4590, NGC 5053, NGC 5904, NGC 6205, NGC 7078, and NGC 7089). At first glance, these findings contradict most models of GC formation that rely on P2 stars being born in a centrally concentrated manner. A possible explanation for this variety of morphologies may, however, also be that the internal dynamical processes (which might play out differently in each GC) are responsible for an early population mixing or inversion of an initially P2-centrally-concentrated GC, to instead show P2 stars in the outer regions.
@@
Generally, the binary–single star scattering is a prevalent cause of stellar ejections from the GC core (see, e.g. Aarseth, 1972; O’Leary et al., 2014; Heggie & Hut, 2003) and is usually mostly visible when the GC goes through core collapse later in its dynamical evolution (see, e.g. Tanikawa et al., 2012, 2013; Fujii & Portegies Zwart, 2014; Pavl´ık & Šubr, 2018). However, suppose a GC contains massive stars initially (O or B-type of several tens of $M_{\odot}$). In that case, they will most likely be paired in binaries (Kroupa, 1995; Goodwin & Kroupa, 2005) and positioned in the core very early on – either the GC is already formed mass segregated, as has been shown for young star clusters (Plunkett et al., 2018; Pavl´ık et al., 2019; Pavl´ık, 2020) but also for older GCs (Zonoozi et al., 2024), or they quickly sink into the core due to mass segregation (e.g. Binney & Tremaine, 2008). Some of those very massive stars would evolve into black holes (BHs) on a time scale of a few Myr (e.g. Hurley et al., 2000, 2002; Belczynski et al., 2008, 2010; Ziosi et al., 2014). If their supernova natal kick velocity is then adequately small (see, e.g. Lyne & Lorimer, 1994; Podsiadlowski et al., 2004), they may be retained in the GC (Peuten et al., 2016; Banerjee, 2017; Baumgardt & Sollima, 2017; Pavl´ık et al., 2018). Hence, BH–BH binaries might be produced promptly in GCs, subsequently ejecting P2 stars from the core to the outer regions through binary–single star scattering.
@@ 2 Theoretical estimates
To achieve mixing (or radial inversion) of P1 and P2, we need to remove a significant fraction of P2 stars (i.e. $N_{\mathrm{encP2}}\sim N_{\rm P2}$) from the core where they were born. Therefore, the time scale for a binary star to encounter and eject most/all of them is $\tau_{\mathrm{enc}}\times N_{\mathrm{encP2}}$.
To observe a mixed/inverted dynamically-young GC, most of the binary–P2-star encounters need to occur early in the cluster evolution, so $\tau_{\mathrm{enc}}\,N_{\mathrm{encP2}}$ per binary must be much lower than the half-mass relaxation time of the cluster (e.g. Spitzer & Hart, 1971)
$\tau_{\mathrm{rh}}(t)\approx\frac{0.138}{\ln{[0.4N(t)]}}\sqrt{\frac{N(t)\,r_{\mathrm{h}}^{3}(t)}{G\,m_{3}}}\,$, (10)
where $N(t)$ is the number of stars in the cluster at time $t$. This is because even if we start from completely P2-centrally-concentrated initial conditions, the P1 stars outside of the core will fall inwards as the cluster relaxes.111 We note that the core relaxation time may be up to an order of magnitude smaller than $\tau_{\mathrm{rh}}$ depending on the stellar density in the core and the velocity distribution (see, e.g. Binney & Tremaine, 2008; Merritt, 2013; Pavl´ık et al., 2024). Thus, we should expect the number of binary–P2-star scatterings to reduce over time and the number of binary–P1-star scatterings to increase. Once the probability of scattering P1 and P2 becomes comparable, the binary would only be able to achieve mixing of the two populations because P1 will get scattered just as much as P2. Consequently, to achieve inversion, the majority of P2 ejections via binary–single scattering need to happen early on and then stop.
@@ 3.1 Initial conditions of the star clusters
We performed a grid of simulations to test the theoretical estimates above. We use clusters with $N=10^{4}$ stars (to probe the parameter space with computational simplicity) and $N=5{\cdot}10^{4}$ stars (for a more realistically sized system). Some of these stars are paired into $N_{\mathrm{bin}}=1$, 3, 10 or 20 massive primordial binaries, and the remaining ones are single equal-mass stars (with $m_{3}=1\,M_{\odot}$). The clusters are isolated, have a Plummer (1911) profile with the half-mass radius $r_{\mathrm{h}}=0.5$, $1.0$, $2.0$ or $4.0\,\mathrm{pc}$, isotropic velocity distribution, and are initialised in virial equilibrium. The initial conditions were generated with McLuster (Küpper et al., 2011) and integrated with the $N$-body tree code PeTar (Wang et al., 2020) in several random realisations each. [^3] Ten realisations of each smaller model and five of the larger models. The simulation time corresponds to several $\tau_{\mathrm{rh}}(0)$ – see also Fig. 1 with the relaxation times of the models for comparison. We note that in Section 5, we compare these models with significantly older GCs in terms of physical units of time (age of several Gyr, whereas our models range between 100 Myr to 2.5 Gyr), however, in terms of their dynamical age (i.e. the number of $\tau_{\mathrm{rh}}$), the observed systems are similar to our models.
@@
We label as ‘P2’ the stars with total energy in the lowest quartile, and the rest are labelled ‘P1’. [^4] We note that the primordial binaries are not counted towards either population, but they do contribute to the potential term in Eq. (11). This separation is based on the assumption that P2 stars are born more concentrated and would have their orbits initially more confined in the central regions. It is usually assumed that P2 would be formed 30–100 Myr after P1 (e.g. Bastian & Lardo, 2018), however, simulating two populations formed at two different times is beyond the scope of this paper. Nonetheless, even our simple distinction between P1 and P2 helps us reveal several effects of the binary–single star interactions and spatial redistribution of the stars.
@@ 3.2 Primordial binary stars
The binary components have $m_{1}=m_{2}=5$ or $30\,M_{\odot}$ depending on the model. [^5] We note that these are typical BH masses and that some detected gravitational-wave sources even exceed $30\,M_{\odot}$. The upper limit of all massive objects per cluster also does not exceed the constraints coming from more realistic systems. Specifically, applying stellar evolution (from Hurley et al., 2000, 2002) on the Kroupa (2001) initial mass function (from $0.08$ to $150\,M_{\odot}$) results in ${\approx}6\,\%$ of the initial total GC mass being contained in BH progenitors. They are generated on circular orbits with semi-major axis $a_{\mathrm{bin}}=5$, 10, 20, or $50\,\mathrm{au}$ – this makes them very hard. [^6] See Eq. (3) where the theoretical hard/soft limit is $a\approx 200\,\mathrm{au}$ for a $(5{+}5)\,M_{\odot}$ binary and ${\approx}8000\,\mathrm{au}$ for a $(30{+}30)\,M_{\odot}$ binary. Even the choice of the widest $50\,\mathrm{au}$ binary is substantially small to theoretically drive the radial redistribution of P1 and P2 stars (see, Fig. 2).
@@
Figure 3: The number of massive binaries in each 10k-* model in time (averaged over each model’s realisations). The models are separated by their initial $r_{\mathrm{h}}$ (columns), primordial binary systems (rows), and their semi-major axes (colours). In each plot, the top horizontal axis shows the evolutionary time in the initial half-mass relaxation times, the lower horizontal axis is in multiples of $\tau_{\mathrm{rh}}$ at the end of the simulation (to mimic the ‘present-day’ value). While we permit the dynamical formation of the low-mass ($m_{3}{+}m_{3}$) or mixed ($m_{1,2}{+}m_{3}$) binaries in the simulations, the numbers here only show the massive ($m_{1}{+}m_{2}$) binaries, including component switching. However, if the primordial binary goes through a state of an unstable triple, it is not counted towards $N_{\mathrm{bin}}$ (which is why we sometimes see the number of binaries decreasing and then increasing again).
@@
Figure 5: The probability density function of the binding energy change, $\alpha$, per interaction, in the 10k-* models. Binary–single and binary–binary interactions are counted together. For reference, we also plot with dashed vertical lines the theoretical values of $\alpha$, for which a single $m_{3}$-star should theoretically escape, as derived using Eq. (8). The figure is separated by each model’s initial half-mass radius (columns), number and masses of the primordial binaries (rows), and the binary semi-major axes (colours).
@@ 4 Evolution of star clusters and binary stars
Fig. 1 further shows that $\tau_{\mathrm{rh}}(t)$ rises over time in the models with more massive binaries. This expansion is greater in the following scenarios: 1) In the most compact clusters, because the relaxation processes are more rapid and the binary–single star interactions can occur more frequently in denser environments (compare the different line styles in each panel of Fig. 1). 2) In the clusters with wider binaries because the scattering cross-section of each binary is larger (e.g. compare the line bifurcation in the models 10k-10x30Mr0.5* in the fourth panel from the left of the figure). 3) In the models with more binaries, as their scattering power increases (compare the same line styles across the panels of the figure).
@@
Consequently, the present-day $\tau_{\mathrm{rh}}$ of the GCs with multiple binaries does not correspond to the relaxation time scale at birth. This means that the dynamical age of such GCs might be underestimated if only inferred from their present-day structure. For comparison, we present some of our results (e.g. in Figs. 4 and 4) in terms of the initial and present-day (i.e. end-of-simulation) $\tau_{\mathrm{rh}}$ simultaneously. And later, when we compare our models with observations, we do so only in terms of the present-day $\tau_{\mathrm{rh}}$.
@@
We must also emphasise that while the dynamical heating increases with the number of primordial binaries in the GC, it is not exactly proportional to it because these binaries also interact with each other and destroy themselves. We show this in Figs. 4 and 4, where we plot the number of massive binaries in our models over time. For instance, see the panels for 10k-30M* and 10k-3x30M* in Fig. 4 – the latter had three binaries initially, however, most of them are quickly destroyed and, on average, only one or two remain after $1\,\tau_{\mathrm{rh}}$, while the former started and finished the simulation with one binary. This binary–binary destruction is more prevalent when the binaries have bigger semi-major axes – see that in most cases, the lighter-coloured lines (tighter binaries) are above the darker-coloured lines (wider binaries) in these figures. Temporary oscillations of $N_{\mathrm{bin}}$ in Figs. 4 and 4 are mainly due to the binary capturing a single star and entering a triple system (occasionally also a quadruple), and we do not count multiples in these figures.
@@
Comparing the plots in Fig. 1 with the highest number of primordial binaries (i.e. 10k-10x30M* and 50k-20x30M*), we see that even with twice as many binaries, $\tau_{\mathrm{rh}}(t)$ does not grow as much in the latter model. This is due to the smaller mass ratio $N_{\mathrm{bin}}(m_{1}{+}m_{2})/M_{\mathrm{cl}}$ in the 50k-20x30M* model than in the 10k-10x30M* model. Moreover, this ratio is further lowered as the number of binary BHs in the 50k-20x30M* model drops to a number between four and ten (see Fig. 4), comparable to the final binary count in 10k-10x30M* (see Fig. 4). A much more numerous BH subsystem or more massive BHs might be necessary to cause an expansion of 50k-20x30M* similar to the 10k-10x30M* model.
@@
In Figs. 6 and 6, we show the distribution of $\alpha$, introduced in Sect. 2 as the change of the binary binding energy per interaction with a single star. We also plot the theoretical lower limit for the interactions that should lead to the escape of the $m_{3}$-stars. Comparing the models of different initial sizes (panels on the same lines in the figures), we see that as $r_{\mathrm{h}}(0)$ of the system increases, the probability of strong interactions leading to escape drops. This is most visible in the models with a single massive binary star (i.e. 10k-30M* in Fig. 6 or 50k-30M* in Fig. 6). While even the wider models with $(30{+}30)\,M_{\odot}$ binaries still produce a population of escaping stars, we may notice that the 10k-5Mr2.0* models with the least massive primordial binary are essentially unable to eject stars (see the top-right panel of Fig. 6), thus any population inversion or mixing in these models should happen at the lowest rate.
@@
While the models with multiple binaries also produce a population of escapers, their distributions of $\alpha$ are more shifted below the limiting value for escape (compare the models in the same column in Figs. 6 or 6, e.g. 50k-*r0.5* in the left-hand panels of Fig. 6). This suggests that as the binary star number in the system increases, a higher number of the single $m_{3}$-stars that pass through the core region are subject to low-energy binary–single interactions rather than a single encounter that would kick them out of the cluster. Consequently, we may expect that the binary–single interactions leading to the mixing (or inversion) of the radial distributions of P1 and P2 would predominantly happen in the more compact clusters with the more massive binary stars. Or in other words, that the more compact clusters would mix more rapidly than the extended clusters of the same mass and binaries.
Figure 8: Same as Fig. 7 but for the 50k-* models.
Figure 10: Same as Fig. 10 but for the 50k-* models.

### v2

@@ 1 Introduction
Star clusters are often used as laboratories for studying star formation, stellar evolution and the dynamics of galaxies. When focusing on the old ($2\,\mathrm{Gyr}$) and relatively massive ($10^{4}\,M_{\odot}$) globular clusters (GCs), these gravitationally bound systems have long been known to host multiple populations of stars with distinct chemical compositions and evolutionary histories (see, e.g. Carretta et al., 2010; Gratton et al., 2012; Bellini et al., 2015; Bastian & Lardo, 2018; Martocchia et al., 2018; Libralato et al., 2019; Tiongco et al., 2019; Martocchia, 2020; Sollima, 2021; Vesperini et al., 2021, and references therein). Hereafter, stars from the primordial (or first) population are referred to as ‘P1’, and the enriched second population as ‘P2’. For instance, variations in the abundances of light elements point to complex formation scenarios of P1 and P2, including the accretion of enriched material onto existing stars or a direct formation of P2 stars out of this enriched material (e.g. Carretta et al., 2010; Bekki, 2011). In some clusters, P1 and P2 are also characterised by different dynamical evolution, which leads to measurable differences in present-day kinematic properties, such as rotation or anisotropy in velocity dispersion (Cordoni et al., 2020b, a; Dalessandro et al., 2024; Leitinger et al., 2025; Cadelano et al., 2024). Since velocity anisotropy affects the relaxation time scales, mass segregation, and the evolution towards energy equipartition (see, e.g. Tiongco et al., 2016; Bianchini et al., 2017; Pavlík & Vesperini, 2021, 2022a, 2022b; Pavlík et al., 2024), it is expected that each stellar population might evolve differently (as shown, e.g. by Livernois et al., 2024).
@@
Most models of GC formation predict that P2 stars are being preferentially born in their cores (e.g. D’Ercole et al., 2008; Lacchin et al., 2022; Yaghoobi et al., 2022a, b). Although the spatial distribution of these populations evolves due to relaxation processes or gas expulsion (e.g. Vesperini et al., 2013; Decressin et al., 2010), dynamically younger GCs are expected to retain their initial structural differences (e.g. Dalessandro et al., 2019, 2024, and references therein). However, a recent observational study by Leitinger et al. (2023) showed that the spatial distribution of P1 and P2 varies among the dynamically-young Galactic GCs. Notably, while P2 stars in some GCs (e.g. NGC 5024 or NGC 6809) appear to be more centrally concentrated than P1 (as expected for GCs less than a few relaxation times old), in other clusters which have similar dynamical ages as these ones (e.g. NGC 3201 or NGC 6101), P2 stars primarily occupy the outer regions of the clusters and P1 is preferentially in the central regions. There are also several dynamically-young GCs which are fully mixed, i.e. they show no preferred concentration of either P1 or P2 (e.g. NGC 288, NGC 4590, NGC 5053, NGC 5904, NGC 6205, NGC 7078, and NGC 7089). At first glance, these findings contradict most models of GC formation that rely on P2 stars being born in a centrally concentrated manner. A possible explanation for this variety of morphologies may, however, also be that the internal dynamical processes (which might play out differently in each GC) are responsible for an early population mixing or inversion of an initially P2-centrally-concentrated GC, to instead show P2 stars in the outer regions.
@@
Generally, the binary–single star scattering is a prevalent cause of stellar ejections from the GC core (see, e.g. Aarseth, 1972; O’Leary et al., 2014; Heggie & Hut, 2003) and is usually mostly visible when the GC goes through core collapse later in its dynamical evolution (see, e.g. Tanikawa et al., 2012, 2013; Fujii & Portegies Zwart, 2014; Pavlík & Šubr, 2018). However, suppose a GC contains massive stars initially (O or B-type of several tens of $M_{\odot}$). In that case, they will most likely be paired in binaries (Kroupa, 1995; Goodwin & Kroupa, 2005) and positioned in the core very early on – either the GC is already formed mass segregated, as has been shown for young star clusters (Plunkett et al., 2018; Pavlík et al., 2019; Pavlík, 2020) but also for older GCs (Zonoozi et al., 2024), or they quickly sink into the core due to mass segregation (e.g. Binney & Tremaine, 2008). Some of those very massive stars would evolve into black holes (BHs) on a time scale of a few Myr (e.g. Hurley et al., 2000, 2002; Belczynski et al., 2008, 2010; Ziosi et al., 2014). If their supernova natal kick velocity is then adequately small (see, e.g. Lyne & Lorimer, 1994; Podsiadlowski et al., 2004), they may be retained in the GC (Peuten et al., 2016; Banerjee, 2017; Baumgardt & Sollima, 2017; Pavlík et al., 2018). Hence, BH–BH binaries might be produced promptly in GCs, subsequently ejecting P2 stars from the core to the outer regions through binary–single star scattering.
@@ 2 Theoretical estimates
To achieve mixing (or radial inversion) of P1 and P2, we need to remove a significant fraction of P2 stars (i.e. $N_{\mathrm{encP2}}\sim N_{\rm P2}$) from the core where they were born. Therefore, the time scale for a binary star to encounter and eject most/all of them is
$\tau_{\mathrm{ej}}\equiv\tau_{\mathrm{enc}}N_{\mathrm{encP2}}\,.$ (10)
However, a single binary can only interact with a fraction of nearby stars whose trajectories are within its cross-section. Thus, to allow most stars in the core to have the opportunity to interact with the binary over $\tau_{\mathrm{ej}}$, their orbits must be sufficiently perturbed by relaxation. This implies that the time scales are $\tau_{\mathrm{rc}}\lesssim\tau_{\mathrm{ej}}$, where $\tau_{\mathrm{rc}}$ is the core relaxation time (see, e.g. Heggie & Hut, 2003), defined as
$\tau_{\mathrm{rc}}(t)=\frac{0.34\,\sigma^{3}_{\rm c}(t)}{\langle m_{\rm c}(t)\rangle\,\rho_{\rm c}(t)\,\ln{[0.4N(t)]}}\,$, (11)
where $\sigma_{\rm c}$ is the core velocity dispersion, $\langle m_{\rm c}\rangle$ is the mean stellar mass in the core, $\rho_{\rm c}$ is the core mass-density, and $N(t)$ is the number of stars in the cluster at time tt. [^1] We note that the core relaxation time may be slightly different depending on the stellar density in the core and the velocity distribution (see, e.g. Binney & Tremaine, 2008; Merritt, 2013; Pavlík et al., 2024).
To observe a mixed/inverted dynamically-young GC, most of the binary–P2-star encounters need to occur early in the cluster evolution, so $\tau_{\mathrm{ej}}$ per binary must be much lower than the half-mass relaxation time of the cluster (e.g. Spitzer & Hart, 1971)
$\tau_{\mathrm{rh}}(t)\approx\frac{0.138}{\ln{[0.4N(t)]}}\sqrt{\frac{N(t)\,r_{\mathrm{h}}^{3}(t)}{G\,m_{3}}}\,.$ (12)
This is because even if we start from completely P2-centrally-concentrated initial conditions, the P1 stars outside of the core will fall inwards as the cluster relaxes. Thus, we should expect the number of binary–P2-star scatterings to reduce over time and the number of binary–P1-star scatterings to increase. Once the probability of scattering P1 and P2 becomes comparable, the binary would only be able to achieve mixing of the two populations because P1 will get scattered just as much as P2. Consequently, to achieve inversion, the majority of P2 ejections via binary–single scattering need to happen early on and then stop.
Figure 3: Evolution of the core relaxation time, $\tau_{\mathrm{rc}}(t)$ in the units of $\tau_{\mathrm{rh}}(t)$, calculated from Eq. (11). The horizontal axis shows time in terms of the initial half-mass relaxation time, $\tau_{\mathrm{rh}}(0)$. The left-hand set of panels is for the 10k-* clusters, the right-hand set is for 50k-*. The models are separated by the primordial binary star numbers and masses (columns), the binary star initial semi-major axes (colours), and the initial cluster half-mass radius (line styles). Each line is averaged over all realisations of the corresponding model.
@@ 3.1 Initial conditions of the star clusters
We performed a grid of simulations to test the theoretical estimates above. We use clusters with $N=10^{4}$ stars (to probe the parameter space with computational simplicity) and $N=5{\cdot}10^{4}$ stars (for a more realistically sized system). Some of these stars are paired into $N_{\mathrm{bin}}=1$, 3, 10 or 20 massive primordial binaries, and the remaining ones are single equal-mass stars (with $m_{3}=1\,M_{\odot}$). The clusters are isolated, have a Plummer (1911) profile with the half-mass radius $r_{\mathrm{h}}=0.5$, $1.0$, $2.0$ or $4.0\,\mathrm{pc}$, isotropic velocity distribution, and are initialised in virial equilibrium. The initial conditions were generated with McLuster (Küpper et al., 2011) and integrated with the $N$-body tree code PeTar (Wang et al., 2020a) in several random realisations each. [^3] Ten realisations of each smaller model and five of the larger models. The simulation time corresponds to several $\tau_{\mathrm{rh}}(0)$ – see also Fig. 3 with the relaxation times of the models for comparison. We note that in Section 5, we compare these models with significantly older GCs in terms of physical units of time (age of several Gyr, whereas our models range between 100 Myr to 2.5 Gyr), however, in terms of their dynamical age (i.e. the number of $\tau_{\mathrm{rh}}$), the observed systems are similar to our models.
@@
We label as ‘P2’ the stars with total energy in the lowest quartile, and the rest are labelled ‘P1’. [^4] We note that the primordial binaries are not counted towards either population, but they do contribute to the potential term in Eq. (13). This separation is based on the assumption that P2 stars are born more concentrated and would have their orbits initially more confined in the central regions. It is usually assumed that P2 would be formed 30–100 Myr after P1 if the cause of enrichment is winds from the asymptotic-giant-branch stars or from interacting binaries (e.g. Bastian & Lardo, 2018). However, in other formation scenarios, where the enrichment is caused by very-/super-massive stars (e.g. Gieles et al., 2018) or stellar-wind mass-loss from stellar merger events (e.g. Wang et al., 2020b), P2 may form within a few Myr after P1. Simulating two populations formed at two different times is beyond the scope of this paper; nevertheless, even our simple distinction between P1 and P2 helps us reveal several effects of the binary–single star interactions and spatial redistribution of the stars.
@@ 3.2 Primordial binary stars
The binary components have $m_{1}=m_{2}=5$ or $30\,M_{\odot}$ depending on the model. [^5] We note that these are typical BH masses and that some detected gravitational-wave sources even exceed $30\,M_{\odot}$. The upper limit of all massive objects per cluster also does not exceed the constraints coming from more realistic systems. Specifically, applying stellar evolution (from Hurley et al., 2000, 2002) on the Kroupa (2001) initial mass function (from $0.08$ to $150\,M_{\odot}$) results in ${\approx}6\,\%$ of the initial total GC mass being contained in BH progenitors. They are generated on circular orbits with semi-major axis $a_{\mathrm{bin}}=5$, 10, 20, or $50\,\mathrm{au}$ – this makes them very hard. [^6] See Eq. (3) where the theoretical hard/soft limit is $a\approx 200\,\mathrm{au}$ for a $(5{+}5)\,M_{\odot}$ binary and ${\approx}8000\,\mathrm{au}$ for a $(30{+}30)\,M_{\odot}$ binary. Even the choice of the widest $50\,\mathrm{au}$ binary is substantially small to theoretically drive the radial redistribution of P1 and P2 stars (see Fig. 1).
@@
Figure 4: The number of massive binaries in each 10k-* model in time (averaged over each model’s realisations). The models are separated by their initial $r_{\mathrm{h}}$ (columns), primordial binary systems (rows), and their semi-major axes (colours). In each plot, the top horizontal axis shows the evolutionary time in the initial half-mass relaxation times, and the lower horizontal axis is in multiples of $\tau_{\mathrm{rh}}$ at the end of the simulation (to mimic the ‘present-day’ value). While we permit the dynamical formation of the low-mass ($m_{3}{+}m_{3}$) or mixed ($m_{1,2}{+}m_{3}$) binaries in the simulations, the numbers here only show the massive ($m_{1}{+}m_{2}$) binaries, including component switching. However, if the primordial binary goes through a state of an unstable triple, it is not counted towards $N_{\mathrm{bin}}$ (which is why we sometimes see the number of binaries decreasing and then increasing again).
@@
Figure 6: The probability density function (PDF) of the binding energy change, $\alpha$, per interaction, in the 10k-* models. Binary–single and binary–binary interactions are counted together. For reference, we also plot with dashed vertical lines the theoretical values of $\alpha$, for which a single $m_{3}$-star should theoretically escape, as derived using Eq. (8). The figure is separated by each model’s initial half-mass radius (columns), number and masses of the primordial binaries (rows), and the binary semi-major axes (colours).
Figure 7: Same as Fig. 7 but for the 50k-* models.
Figure 9: Same as Fig. 8 but for the 50k-* models.
@@ 4 Evolution of star clusters and binary stars
Similarly, we can infer from Fig. 3 that the cores of the clusters with a single low-mass binary (i.e. 10k-5M* and 50k-30M* models) gradually proceed towards core collapse. On the other hand, the cores of the models with massive or multiple binary stars contract more rapidly initially. In the case of the models 50k-10x30M*, 50k-20x30M*, and to an extent also 10k-10x30M, Fig. 3 also shows an increase of $\tau_{\mathrm{rc}}$ after the initial drop, corresponding to the expansion of the core due to binary heating.
Fig. 3 further shows that $\tau_{\mathrm{rh}}(t)$ rises over time in the models with more massive binaries. This expansion is greater in the following scenarios:
  1. 1.
@@
    In the most compact clusters, because the relaxation processes are more rapid and the binary–single star interactions can occur more frequently in denser environments (compare the different line styles in each panel of Fig. 3).
  2. 2.
    In the clusters with wider binaries, because the scattering cross-section of each binary is larger (e.g. compare the line bifurcation in the models 10k-10x30Mr0.5* in the fourth panel from the left of the figure).
  3. 3.
    In the models with more binaries, as their scattering power increases (compare the same line styles across the panels of the figure).
@@
Consequently, the present-day $\tau_{\mathrm{rh}}$ of the GCs with multiple binaries does not correspond to the relaxation time scale at birth. This means that the dynamical age of such GCs might be underestimated if only inferred from their present-day structure. For comparison, we present some of our results (e.g. in Figs. 5 and 5) in terms of the initial and present-day (i.e. end-of-simulation) $\tau_{\mathrm{rh}}$ simultaneously. And later, when we compare our models with observations, we do so only in terms of the present-day $\tau_{\mathrm{rh}}$.
@@
We must also emphasise that while the dynamical heating increases with the number of primordial binaries in the GC, it is not exactly proportional to it because these binaries also interact with each other and destroy themselves. We show this in Figs. 5 and 5, where we plot the number of massive binaries in our models over time. For instance, see the panels for 10k-30M* and 10k-3x30M* in Fig. 5 – the latter had three binaries initially, however, most of them are quickly destroyed and, on average, only one or two remain after $1\,\tau_{\mathrm{rh}}$, while the former started and finished the simulation with one binary. This binary–binary destruction is more prevalent when the binaries have larger semi-major axes – see that in most cases, the lighter-coloured lines (tighter binaries) are above the darker-coloured lines (wider binaries) in these figures. Temporary oscillations of $N_{\mathrm{bin}}$ in Figs. 5 and 5 are mainly due to the binary capturing a single star and entering a triple system (occasionally also a quadruple), and we do not count multiples in these figures.
@@
Comparing the plots in Fig. 3 with the highest number of primordial binaries (i.e. 10k-10x30M* and 50k-20x30M*), we see that even with twice as many binaries, $\tau_{\mathrm{rh}}(t)$ does not grow as much in the latter model. This is due to the smaller mass ratio $N_{\mathrm{bin}}(m_{1}{+}m_{2})/M_{\mathrm{cl}}$ in the 50k-20x30M* models than in the 10k-10x30M* models. Moreover, this ratio is further lowered as the number of binary BHs in the 50k-20x30M* models drops to a number between four and ten (see Fig. 5), comparable to the final binary count in 10k-10x30M* (see Fig. 5). A much more numerous BH subsystem or more massive BHs might be necessary to cause an expansion of 50k-20x30M* similar to the 10k-10x30M* model.
Moreover, Fig. 3 further shows that the ratio $\tau_{\mathrm{rc}}/\tau_{\mathrm{rh}}$ in the 50k-20x30M* models is higher than in the 10k-10x30M* models. This signals (as discussed in Sec. 2) that the binaries in the 50k-20x30M* models cannot interact with as many core P2 stars as in the 10k-10x30M* models, and the population mixing in the former models would be less effective.
@@
In Figs. 7 and 7, we show the distribution of $\alpha$, introduced in Sect. 2 as the change of the binary binding energy per interaction with a single star. We also plot the theoretical lower limit for the interactions that should lead to the escape of the $m_{3}$-stars. Comparing the models of different initial sizes (panels on the same lines in the figures), we see that as $r_{\mathrm{h}}(0)$ of the system increases, the probability of strong interactions leading to escape drops. This is most visible in the models with a single massive binary star (i.e. 10k-30M* in Fig. 7 or 50k-30M* in Fig. 7). While even the wider models with $(30{+}30)\,M_{\odot}$ binaries still produce a population of escaping stars, we may notice that the 10k-5Mr2.0* models with the least massive primordial binary are essentially unable to eject stars (see the top-right panel of Fig. 7), thus any population inversion or mixing in these models should happen at the lowest rate.
@@
While the models with multiple binaries also produce a population of escapers, their distributions of $\alpha$ are more shifted below the limiting value for escape (compare the models in the same column in Figs. 7 or 7, e.g. 50k-*r0.5* in the left-hand panels of Fig. 7). This suggests that as the binary star number in the system increases, a higher number of the single $m_{3}$-stars that pass through the core region are subject to low-energy binary–single interactions rather than a single encounter that would kick them out of the cluster. Consequently, we may expect that the binary–single interactions leading to the mixing (or inversion) of the radial distributions of P1 and P2 would predominantly happen in the more compact clusters with the more massive binary stars. Or in other words, that the more compact clusters would mix more rapidly than the extended clusters of the same mass and binaries.